# PointMapPolicy: Structured Point Cloud Processing for Multi-Modal Imitation Learning

### v2

@@ 4.2 Point Cloud Encoding
We conducted controlled experiments on RoboCasa using identical xLSTM backbones with different point cloud processing encoders: 1) PointNet-xyz: Following DP3 [ze20243d], we gather point clouds from 3 camera views and use Furthest Point Sampling (FPS) to downsample to 1024 points, then apply MLP with maxpooling to create a compact 3D token. 2) PointNet-color: Same process as PointNet-xyz but using colored points with XYZRGB information. 3) PointPatch: We use FPS to sample 256 center points, apply k-Nearest Neighbors to create 256 point patches with 32 points each, tokenize each patch using MLP with maxpooling, then process the resulting tokens with a transformer to generate compact 3D representations. 4) 3D-Lifting: We extract CLIP features (frozen) from each camera view and lift the 2D features into 3D space, then use a transformer to process the lifted tokens. The 3D tokens are then passed to the diffusion policy with an identical X-Block backbone.
@@ 4.3 Ablation Study
Understanding Model Attention Patterns. To uncover where the model attends during action prediction, we apply Grad-CAM++ [grad_cam_plus_plus] to highlight the regions most influential for action decisions across different modalities. For detailed visualizations, see Appendix D.
@@ 5.3 Real-World Main Results
Tasks
Demos
Per Task
Methods with Scores
@@ 8 Acknowledgments
This work was supported by funding from the German Research Foundation (DFG, German Research Foundation) - SFB-1574 – 471687386, and is part of the DFG AI Research Unit 5339. Additional funding was provided by the pilot program Core Informatics of the Helmholtz Association (HGF). NS and GN were supported by the Carl Zeiss Foundation under the project JuBot (Jung Bleiben mit Robotern). Xiaogang Jia and Xinkai Jiang acknowledge the support from the China Scholarship Council (CSC). This work has also received funding from the European Union’s Horizon Europe research and innovation programme under Grant Agreement No. 101171393 (SMARTI³ – "Scalable Manipulation Learning through AR-enhanced Teleoperation enabling Intuitive Interactive Instructions"). The authors also acknowledge support by the state of Baden-Württemberg through the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg, and by the German Federal Ministry of Education and Research.
@@ A.1 RoboCasa Benchmark
RoboCasa [nasiriany2024robocasa] is a large-scale simulation framework developed to train generalist robots in realistic and diverse home environments, with a particular focus on kitchen scenarios. The benchmark comprises 100 tasks, including 25 atomic tasks with 50 human demonstrations and 75 composite tasks with auto-generated demonstrations. These tasks are centered around eight fundamental robotic skills relevant to real-world home environments: (1) pick-and-place, (2) opening and closing doors, (3) opening and closing drawers, (4) twisting knobs, (5) turning levers, (6) pressing buttons, (7) insertion, and (8) navigation.
@@ A.2 CALVIN Benchmark
The CALVIN benchmark [mees2022calvin] is a long-horizon manipulation benchmark featuring four visually distinct tabletop environments (A–D), each containing a common set of objects and 34 manipulation tasks. Agents are given natural language instructions describing sequences of up to 5 tasks to be executed in order. The primary evaluation involves completing 1000 such instruction chains in environment D. Agents are scored by the number of tasks successfully completed per chain, with a maximum rollout length of 5.
@@
  • HULC [mees2022hulc]: a hierarchical imitation learning framework that combines a discrete skill planner (learned via VAE) with a low-level controller trained on dense teleoperation data. It does not use language pretraining or frozen encoders.
  • LAD [zhang2022lad]: extends HULC by replacing the planner with a high-level diffusion model that learns to generate latent skills conditioned on goals. The policy uses a U-Net-based diffusion process and CLIP-encoded language inputs.
  • Distill-D [ha2023scaling]: a CLIP-conditioned continuous-time diffusion policy. It distills robot behaviors from demonstrations using dense supervision and large batch sizes, incorporating CLIP goals and continuous action prediction.
  • MT-ACT [liu2023libero]: a transformer-based policy that predicts temporally extended action chunks. It encodes entire trajectories using VAE-style representation learning and uses a transformer encoder-decoder for action generation.
  • RoboFlamingo [li2023vision]: based on OpenFlamingo, this model integrates a frozen VLM with a lightweight policy head. It is pretrained on large-scale vision-language data and finetuned on CALVIN using supervised behavior cloning.
  • SuSIE [wu2024susie]: a scalable instruction-following diffusion policy. It is pre-trained on curated robot demonstrations and uses an instruction-conditioned denoising process with significant offline finetuning.
  • GR-1 [wu2023unleashing]: a powerful decoder-only transformer trained on large-scale synthetic video data. The model is capable of generating long sequences of actions and is finetuned on CALVIN for grounding.
  • CLOVER [bu2024clover]: a video diffusion planner that predicts intermediate visual goals via video generation and closes the loop using low-level policy feedback. It does not require internet-scale pretraining and achieves strong multi-step rollout success.
  • MoDE [zhou2024mode]: Mixture-of-Diffusion-Experts model with sparse routing. It supports both small (non-pretrained) and large (pretrained) variants. The pretrained variant achieves top performance while maintaining low inference cost.
  • Seer / Seer-Large [zhou2024seer]: large-scale transformer models pretrained on 1000+ hours of robot play data. Seer incorporates language, vision, and action streams into a unified transformer and achieves strong generalization, particularly when scaled up.
Results.
[…]
A.3 CALVIN Evaluation Details
CALVIN Task Suite.
The CALVIN benchmark [mees2022calvin] consists of four environments (A–D) with identical object sets and tasks but differing in visual appearance and lighting. Each episode comprises natural language instructions describing a chain of 5 atomic manipulation goals selected from 34 total tasks. These include pushing, pulling, opening drawers, pressing buttons, etc. The robot receives RGB-D observations from multiple camera views and proprioceptive information. The action space is 7-dimensional, including end-effector translation, rotation, and gripper control.
Language-Conditioned Evaluation.
Agents are evaluated on their ability to execute long-horizon natural language commands consisting of 5 consecutive tasks. The evaluation split uses 1000 unique instruction chains in environment D. The model receives a reward of 1 upon successful completion of each task in the chain, and evaluation metrics report the average number of tasks completed per rollout over 3 random seeds.
[…]
To evaluate generalization, the model is trained on environments A, B, and C, and tested zero-shot on environment D. No data from environment D is used during training. This split emphasizes robustness to domain shifts and visual changes.
@@ Appendix D Activation Map Analysis
To gain qualitative insights into what regions the visual encoders attend to during action inference, we visualize activation maps using Grad-CAM++ [grad_cam_plus_plus]. Unlike classification tasks, our diffusion-based policy does not predict discrete categories, therefore, we apply Grad-CAM++ using the diffusion loss as the target signal, following the approach of highlighting input regions that most influence the denoised trajectory prediction. We generate the heatmaps using the Grad-CAM++ implementation [^1] https://github.com/jacobgil/pytorch-grad-cam, and compute activations for each camera view across three RoboCasa tasks: OpenDrawer, Turn On Sink Faucet, and Coffee Serve Mug. In all figures, we use a ConvNeXtv2 encoder and extract Grad-CAM++ heatmaps from the final convolutional block before normalization. Each visualization consists of six images arranged in two rows. The top row shows the raw visual input (RGB or XYZ visualized in color), and the bottom row displays the corresponding Grad-CAM++ heatmaps for each of the three camera views: static left, static right, and wrist-mounted. These maps highlight spatial regions with the greatest impact on predicted actions.

### v3

@@ 4.2 Point Cloud Encoding
We conducted controlled experiments on RoboCasa using identical xLSTM backbones with different point cloud processing encoders: 1) PointNet-xyz: Following DP3 [50], we gather point clouds from 3 camera views and use Furthest Point Sampling (FPS) to downsample to 1024 points, then apply MLP with maxpooling to create a compact 3D token. 2) PointNet-color: Same process as PointNet-xyz but using colored points with XYZRGB information. 3) PointPatch: We use FPS to sample 256 center points, apply k-Nearest Neighbors to create 256 point patches with 32 points each, tokenize each patch using MLP with maxpooling, then process the resulting tokens with a transformer to generate compact 3D representations. 4) 3D-Lifting: We extract CLIP features (frozen) from each camera view and lift the 2D features into 3D space, then use a transformer to process the lifted tokens. The 3D tokens are then passed to the diffusion policy with an identical X-Block backbone.
@@ 4.3 Ablation Study
Understanding Model Attention Patterns. To uncover where the model attends during action prediction, we apply Grad-CAM++ [8] to highlight the regions most influential for action decisions across different modalities. For detailed visualizations, see Appendix D.
@@ 5.3 Real-World Main Results
Tasks Demos Per Task Methods with Scores
@@ 8 Acknowledgments
This work was supported by the European Research Council (ERC) under the European Union’s Horizon Europe programme through the project SMARTI³ (Grant Agreement No. 101171393). The authors also acknowledge funding from the German Research Foundation (DFG) within the framework of Collaborative Research Centre SFB 1574 “Circular Factory for the Perpetual Product” (Project No. 471687386). This work was also supported by funding from the pilot program Core Informatics of the Helmholtz Association (HGF). NS and GN were supported by the Carl Zeiss Foundation under the project JuBot (Jung Bleiben mit Robotern). Xiaogang Jia and Xinkai Jiang acknowledge the support from the China Scholarship Council (CSC). The authors also acknowledge support by the state of Baden-Württemberg through the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg, and by the German Federal Ministry of Education and Research.
@@ A.1 RoboCasa Benchmark
RoboCasa [32] is a large-scale simulation framework developed to train generalist robots in realistic and diverse home environments, with a particular focus on kitchen scenarios. The benchmark comprises 100 tasks, including 25 atomic tasks with 50 human demonstrations and 75 composite tasks with auto-generated demonstrations. These tasks are centered around eight fundamental robotic skills relevant to real-world home environments: (1) pick-and-place, (2) opening and closing doors, (3) opening and closing drawers, (4) twisting knobs, (5) turning levers, (6) pressing buttons, (7) insertion, and (8) navigation.
@@ A.2 CALVIN Benchmark
The CALVIN benchmark [31] is a long-horizon manipulation benchmark featuring four visually distinct tabletop environments (A–D), each containing a common set of objects and 34 manipulation tasks. Agents are given natural language instructions describing sequences of up to 5 tasks to be executed in order. The primary evaluation involves completing 1000 such instruction chains in environment D. Agents are scored by the number of tasks successfully completed per chain, with a maximum rollout length of 5.
@@
  • RoboFlamingo [27]: based on OpenFlamingo, this model integrates a frozen VLM with a lightweight policy head. It is pretrained on large-scale vision-language data and finetuned on CALVIN using supervised behavior cloning.
  • SuSIE [5]: a scalable instruction-following diffusion policy. It is pre-trained on curated robot demonstrations and uses an instruction-conditioned denoising process with significant offline finetuning.
  • GR-1 [48]: a powerful decoder-only transformer trained on large-scale synthetic video data. The model is capable of generating long sequences of actions and is finetuned on CALVIN for grounding.
  • CLOVER [7]: a video diffusion planner that predicts intermediate visual goals via video generation and closes the loop using low-level policy feedback. It does not require internet-scale pretraining and achieves strong multi-step rollout success.
  • MoDE [40]: Mixture-of-Diffusion-Experts model with sparse routing. It supports both small (non-pretrained) and large (pretrained) variants. The pretrained variant achieves top performance while maintaining low inference cost.
  • Seer / Seer-Large [45]: large-scale transformer models pretrained on 1000+ hours of robot play data. Seer incorporates language, vision, and action streams into a unified transformer and achieves strong generalization, particularly when scaled up.
@@ Appendix D Activation Map Analysis
To gain qualitative insights into what regions the visual encoders attend to during action inference, we visualize activation maps using Grad-CAM++ [8]. Unlike classification tasks, our diffusion-based policy does not predict discrete categories, therefore, we apply Grad-CAM++ using the diffusion loss as the target signal, following the approach of highlighting input regions that most influence the denoised trajectory prediction. We generate the heatmaps using the Grad-CAM++ implementation [^1] https://github.com/jacobgil/pytorch-grad-cam, and compute activations for each camera view across three RoboCasa tasks: OpenDrawer, Turn On Sink Faucet, and Coffee Serve Mug. In all figures, we use a ConvNeXtv2 encoder and extract Grad-CAM++ heatmaps from the final convolutional block before normalization. Each visualization consists of six images arranged in two rows. The top row shows the raw visual input (RGB or XYZ visualized in color), and the bottom row displays the corresponding Grad-CAM++ heatmaps for each of the three camera views: static left, static right, and wrist-mounted. These maps highlight spatial regions with the greatest impact on predicted actions.